OpenStudy (anonymous):

In order to start a small business a student takes out a simple interest loan for $7000 for 9 months with 12% interest. How much interest must the student pay (round to the nearest cent) and find the future value of the loan (round to the nearest cent)

OpenStudy (anonymous):

@texaschic101

OpenStudy (danjs):

REmember the definition for the simple interest Simple Interest = Current Value x rate x time in years I = P*r*t just have to calculate the interest I

OpenStudy (danjs):

r = 0.12 P = 7000 t = 9/12

OpenStudy (texaschic101):

but what about the future value ? Thats where I get lost

OpenStudy (danjs):

yeh , i guess they want that in terms of time maybe,

OpenStudy (texaschic101):

does this have anything to do with it...or am I way off... http://www.algebra.com/algebra/homework/Finance/PMTFV.lesson

OpenStudy (anonymous):

I'm not sure

OpenStudy (danjs):

Future Value = Start Value + Accumulated interest

OpenStudy (texaschic101):

so basically do the I = PRT to find I, and when that is found, add it back to P ??

OpenStudy (danjs):

value of the loan at 9months is the future i assume, not t

OpenStudy (danjs):

FV = P + P*r*t FV = P(1+r*t)

OpenStudy (texaschic101):

oh....ok...I see now. FV = 7000 + 7000 * 0.12 * 9/12(or 3/4)

OpenStudy (texaschic101):

you get it smiley ?

OpenStudy (anonymous):

kinda

OpenStudy (texaschic101):

FV = 7000 + 7000 * 0.12 * 3/4 FV = 7000 + 840 * 3/4 FV = 7000 + 630 FV = 7630 According to the formula that Dan gave us....I got this

OpenStudy (anonymous):

oh alright what was the amount of interest